# A Novel Hierarchical Co-Optimization Framework for Coordinated Task Scheduling and Power Dispatch in Computing Power Networks

### v1

@@ IV-C DRL-based Real-Time CPN Task Scheduling
      CPN State: Characteristics of tasks in the queue (e.g., resource requirements, deadlines), current resource utilization and power consumption of each CPN node.
@@
  • Action ($a_{t}\in\mathcal{A}$): For the task at the head of the queue, the agent selects an action from a discrete set. An action is a tuple $(n,\tau_{type})$ representing the decision to assign the task to node $n$ to be processed by resource type $\tau_{type}$ (e.g., CPU, GPU). The action space also includes deferring the task.
@@
    The reward is then penalized by an additional term proportional to $Q_{t}\times\text{Carbon}_{t}$, which strongly discourages carbon-intensive actions when the system is already over its carbon budget. And the $\text{Penalty}_{t}$ represents a large negative penalty for missing a task’s deadline.
@@ 5. Obtain real-time LMPs and MCIs for all CPN node locations
   6. Construct state vector $s_{t}\leftarrow$ (CPN state + Grid state)
   7. DRL agent takes action $a_{t}\leftarrow\pi(s_{t})$ to schedule next CPN task
@@ 8. Update CPN power demand $P_{i,t}^{CPN}$ based on action $a_{t}$
   9. Execute scheduled task and power dispatch
   10. Update BESS SOC and CPN resource status
@@ IV-D1 Stage 1: Day-Ahead SUC
  • Subproblems: For each of the $|\Omega|$ scenarios, a linear program (LP) is solved. The complexity of solving an LP with modern interior-point methods is polynomial in the number of variables and constraints. Since the subproblems are independent for a given commitment schedule, they can be solved in parallel. The time taken per iteration for this step is thus equivalent to solving a single LP.
@@ V Performance Evaluation
This section presents a comprehensive empirical validation of the proposed Two-Stage Co-Optimization (TSCO) framework. A high-fidelity simulation environment is developed to assess its performance in terms of economic efficiency, environmental impact, grid stability, and CPN QoS.
@@ V-A Simulation Setup
To ground our environmental and economic calculations in reality, we incorporate two external data sources. Real-time marginal carbon intensity (MCI) data is obtained via the WattTime API, which provides 5-minute resolution data on the emissions impact of consuming an additional MWh of electricity in various grid regions [^8] https://watttime.org/data-science/data-signals/marginal-co2/. Historical hourly locational marginal price (LMP) data is sourced from the California Independent System Operator (CAISO) public database [^9] https://www.gridstatus.io/live/caiso.
@@ V-B Comparison Schemes
  • Cost-Only Optimizer (CO-Opt) [32]: This baseline uses the same two-stage optimization architecture as TSCO but with the carbon price set to zero. This represents the current industry-standard approach of economic dispatch, which focuses exclusively on minimizing direct economic costs without explicit consideration for environmental impact.
@@
  • Decoupled Framework (DC-Frame) [34]: This baseline represents the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately. The power system operation is optimized first to generate a fixed 24-hour profile of electricity prices and carbon intensities. Subsequently, the CPN scheduler optimizes its task scheduling based on these static, pre-computed signals. This approach is carbon-aware but lacks the tight, real-time feedback loop of our co-optimization framework.
V-C Simulation Results
@@ V-C1 Baseline Performance Comparison
In this scenario, all four methods were simulated over one week (168 hours) with the carbon price set at a representative value of \$50\\$50/ton. Each data point presented is the result of $10$ independent simulation runs, from which we calculate the mean and standard deviation and present in Fig. 3.
@@ V-C2 Sensitivity to Carbon Price
To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from \$25\\$25/ton to \$150\\$150/ton. This analysis focuses on the TSCO and DC-Frame methods, as the other two baselines are insensitive to carbon price by design. The comprehensive results are presented in Fig. 4.
The TSCO framework consistently achieves a better trade-off. For any given carbon emission level, it operates at a lower cost than the DC-Frame. As the carbon price increases, both methods are incentivized to reduce emissions, but TSCO does so more efficiently. This superior performance is a direct result of the real-time feedback loop between the grid and the CPN, which allows the system to find more efficient operating points. The RG-Sched method is shown as fixed points, as their operational strategy does not adapt to the carbon price.
@@ VI Conclusion
The findings of this research have significant practical implications for both power system operators and the owners of large-scale computational infrastructure. For grid operators, flexible CPNs can serve as a valuable source of ancillary services, reducing the need for expensive battery storage or fossil-fuel reserves. For CPN operators, participating in such programs can create new revenue streams and drastically reduce operational expenditures and carbon footprints.

### v2

@@ IV-C DRL-based Real-Time CPN Task Scheduling
      CPN State: Characteristics of tasks in the queue (e.g., resource requirements, deadlines), current resource and bandwidth utilization, and power consumption of each CPN node.
@@
    Furthermore, the state vector $s_{t}$ is augmented with two day-ahead constraint tracking metrics:
    $Remaining_{n,h,t}=C_{n,h,t}^{comp,reserve}-\sum_{k,\tau}r_{\tau,h}x_{k,\tau,n,h,t-1},$ (23)
    $Budget_{t}=E_{budget}-\sum_{\tau=1}^{t}\sum_{\omega\in\Omega}\pi_{\omega}E_{\tau,\omega},$ (24)
    The former remains a computational resource reserve of node $n$ at time $t$. The latter remains the carbon budget for the planning horizon. These metrics enable the DRL agent to proactively avoid approaching constraint boundaries.
  • Action ($a_{t}\in\mathcal{A}$): For the task at the head of the queue, the agent selects an action from a discrete set. An action is a tuple $(n,h,\tau_{type})$ representing the decision to assign the task to node $n$ to be processed by resource type $\tau_{type}$ (e.g., CPU, GPU). The action space also includes deferring the task. The action space is further constrained by the day-ahead resource reservations from Stage 1. For any task assignment action $(n,h,\tau_{type})$, the total resource demand of tasks scheduled on node $n$ at time $t$, i.e., $\sum_{k,\tau}r_{\tau,h}x_{k,\tau,n,h,t}$ must not exceed the day-ahead reserved capacity $C_{n,h,t}^{comp,reserve}$. This constraint is enforced directly in the DRL agent’s action selection process by masking infeasible assignments, e.g., nodes with remaining capacity $\leq$ task resource requirement $r_{\tau},h$.
@@
    The reward is then penalized by an additional term proportional to $Q_{t}\text{Carbon}_{t}$, which strongly discourages carbon-intensive actions when the system is already over its carbon budget. And the $\text{Penalty}_{t}$ represents a large negative penalty for missing a task’s deadline. The intuition behind this virtual carbon queue is to track the deviation between cumulative carbon emissions and the long-term budget. $Q_{t}$ increases when the current carbon emission $\text{Carbon}_{t}$ exceeds the average allowable emission $E_{budget}/T$ and resets to 0 if emissions are within the budget. When $Q_{t}$ grows, indicating the system is approaching or exceeding the carbon budget, the additional penalty term $Q_{t}\text{Carbon}_{t}$ in the reward function strongly discourages the DRL agent from assigning tasks to high-carbon-intensity nodes. It then ensures that short-term scheduling decisions do not violate the long-term sustainability goal while avoiding excessive sacrifice of economic efficiency.
  • Constraint Adherence and Feedback Mechanism: To address potential deviations from the day-ahead plan, a two-tier feedback mechanism is implemented:
    1. 1.
      Real-Time Infeasibility Correction: If the real-time ED detects that the CPN power demand $P_{i,t}^{CPN}$ violates grid constraints, such as transmission line limits or generator ramp constraints, the ED module sends a “constraint violation signal” to the DRL agent. The agent then temporarily narrows its action space by increasing the penalty weight $w_{pen}$ for tasks assigned to nodes causing violations, or masking those nodes for 1-2 time steps until the grid state stabilizes.
    2. 2.
      Rolling Day-Ahead Plan Update: If the cumulative deviation between real-time CPN resource utilization and day-ahead reserve, i.e., $\sum_{t}|\sum_{k,\tau}r_{\tau}x_{k,\tau,n,t}-C_{n,t}^{comp,reserve}|$, exceeds a predefined threshold over 4 consecutive hours, the day-ahead SUC module is triggered to perform a mid-day rolling update. Then, the algorithm adjusts the remaining 24-hour reserve allocation and generator commitment status based on the latest real-time data.
@@ 5. Obtain real-time LMPs and MCIs for all CPN node locations
   6. Construct state vector $s_{t}\leftarrow$ (CPN state + Grid state + resource/carbon budget)
   7. Mask infeasible actions, DRL agent takes action $a_{t}\leftarrow\pi(s_{t})$ to schedule task
@@ 8. Update CPN power demand $P_{i,t}^{CPN}$ based on action $a_{t}$
   9. Correct grid constraint violations (if any), execute dispatch/scheduling
   10. Update BESS SOC, CPN resource status and virtual carbon queue
@@ IV-D1 Stage 1: Day-Ahead SUC
  • Subproblems: For each of the $\Omega$ scenarios, a linear program (LP) is solved. The complexity of solving an LP with modern interior-point methods is polynomial in the number of variables and constraints. Since the subproblems are independent for a given commitment schedule, they can be solved in parallel. The time taken per iteration for this step is thus equivalent to solving a single LP.
@@ V Performance Evaluation
This section presents a comprehensive empirical validation of the proposed Two-Stage Co-Optimization (TSCO) framework. A high-fidelity simulation environment is developed to assess its performance in terms of economic efficiency, environmental impact, grid stability, and CPN QoS. Additional experiments on convergence, scalability, ablation, and computational burden are conducted to further verify the framework’s engineering feasibility and core value.
@@ V-A Simulation Setup
To ground our environmental and economic calculations in reality, we incorporate two external data sources. Real-time marginal carbon intensity (MCI) data is obtained via the WattTime API, which provides 5-minute resolution data on the emissions impact of consuming an additional MWh of electricity in various grid regions [^8] https://watttime.org/data-science/data-signals/marginal-co2/. Historical hourly locational marginal price (LMP) data is sourced from the California Independent System Operator (CAISO) public database [^9] https://www.gridstatus.io/live/caiso. To ensure the generality of the experimental results, all values are presented as the mean ± standard deviation derived from 10 independent runs.
@@ V-B Comparison Schemes
  • Cost-Only Optimizer (CO-Opt) [35]: This baseline uses the same two-stage optimization architecture as TSCO but with the carbon price set to zero. It represents the current industry-standard approach of economic dispatch, which focuses exclusively on minimizing direct economic costs without consideration for environmental impact.
@@
  • Decoupled Framework (DC-Frame) [37]: It is the state-of-the-art in carbon-aware computing, where the power system and CPN are optimized separately. The power system operation is optimized first to generate a fixed 24-hour profile of electricity prices and carbon intensities. Subsequently, the CPN scheduler optimizes its task scheduling based on these static, pre-computed signals. This approach is carbon-aware but lacks the tight, real-time feedback loop of our co-optimization framework.
V-C Comparison Performances
@@ V-C1 Baseline Performance Comparison
In this scenario, all four methods were simulated over one week (168 hours) with the carbon price set at a representative value of $\mathdollar 50$/ton. The mean and standard deviation are present in Fig. 3.
@@ V-C2 Sensitivity to Carbon Price
To analyze the trade-off between economic and environmental objectives, we varied the carbon price $\lambda_{CO_{2}}$ from $\mathdollar 25$/ton to $\mathdollar 150$/ton. This analysis focuses on the TSCO and DC-Frame methods, as the other two baselines are insensitive to carbon price by design. The comprehensive results are presented in Fig. 4.
The TSCO framework consistently achieves a better trade-off. For any given carbon emission level, it operates at a lower cost than the DC-Frame. As the carbon price increases, both methods are incentivized to reduce emissions, but TSCO does so more efficiently. This superior performance is a direct result of the real-time feedback loop between the grid and the CPN, which allows the system to find more efficient operating points. The RG-Sched method is shown as a fixed point, as their operational strategy does not adapt to the carbon price.
TABLE II: Convergence Performance Comparison
Method Load Intensity Convergence Iterations Convergence Time (min) Bound Gap (\$)
Direct-MILP Low - (infeasible) Larger than 360 (timeout) -
Medium - (infeasible) Larger than 360 (timeout) -
High - (infeasible) Larger than 360 (timeout) -
Improved-Benders Low 28.3±2.1 45.7±3.2 128.5±15.3
Medium 35.6±2.8 58.2±4.1 156.3±18.7
High 42.1±3.5 72.4±5.6 189.7±21.2
TSCO-Benders Low 21.5±1.8 32.4±2.5 98.6±12.4
Medium 27.8±2.3 43.1±3.7 121.4±16.5
High 33.2±2.9 55.8±4.8 145.2±19.3
@@
V-D Convergence Comparison
To verify the efficiency and stability of the Benders decomposition in the day-ahead SUC stage, we compare two methods under different load intensities, including Direct-MILP without decomposition and Improved-Benders in [32]. The results are summarized in Table II. Among them, the mid-range workload represents an average daily task arrival rate of $800$ jobs per day with a total computational workload of $5\times 10^{15}$ FLOPs, and an average resource requirement of $4$ CPU cores and 2 GPU cores or equivalent TPU resources per job. And the low and high loads are $0.8$ times and $1.2$ times the medium load, respectively.
Direct-MILP fails to converge within 6 hours for all scenarios due to the high dimensionality of the SUC problem. Both Benders-based methods achieve convergence, but TSCO-Benders outperforms Improved-Benders by $24.2\%-27.9\%$ in convergence iterations and $29.1\%-31.3\%$ in convergence time. This is because the tailored Benders cuts in TSCO effectively reduce the search space of the master problem. The bound gap of TSCO-Benders is consistently below $\mathdollar 150$, indicating high solution accuracy. For medium load, TSCO-Benders takes an average of $43.1$ minutes to converge, which is acceptable for day-ahead planning.
V-E Scalability Verification
We also expand the system scale in three dimensions to test the scalability of the TSCO framework. The results are shown in Table III.
TABLE III: Scalability Performance Under Different System Scales
CPN Nodes RES Scenarios Gen. Count Day-Ahead Conv. Time (min) Day-Ahead Conv. Iterations Real-Time ED Time (s/step) DRL Inference Time (ms/task) Performance Retention Rate (%)
5 100 6 43.1±3.7 27.8±2.3 1.2±0.1 8.5±0.6 100.0±0.0
10 200 8 68.5±4.9 35.2±2.7 1.8±0.2 10.3±0.8 96.7±0.8
15 300 10 95.3±6.2 42.7±3.1 2.5±0.3 12.1±1.0 93.5±1.1
20 500 12 132.7±7.8 51.4±3.6 3.3±0.4 14.8±1.2 90.2±1.5
Figure 5: Ablation test. (a) Total operational cost; (b) Total carbon emissions; (c) RES curtailment; (d) CPN job success rate; (d) Average job tardiness.
As the system scale expands with 20 nodes, 500 scenarios, and 12 generators, the day-ahead convergence time increases to 132.7 minutes. It is still within the 24-hour planning window, and the real-time ED/DRL time remains at the millisecond-to-second level. The performance retention rate is $90.2\%$, meaning the carbon emission reduction rate and RES utilization only decrease by $9.8\%$ compared to the baseline. This demonstrates that TSCO’s hierarchical structure effectively isolates computational complexity. The day-ahead stage tolerates longer computation for large-scale optimization, while the real-time stage maintains fast response via LP and DRL inference. The framework thus meets the requirements of large-scale CPN-grid integration.
V-F Ablation Test
To validate the necessity of key modules in the TSCO framework, we design four variants by removing core components one by one. In this test, we remove the nonlinear power consumption model of CPN from the TSCO and replace it with a simple linear model; Constraints of the communication link; Carbon-aware DRL; and integration with the two-stage process in TSCO. Its simulation parameters are set in accordance with those of Section V-C-1).
The results are presented in Fig. 5. Specifically, the nonlinear power model reduces carbon emissions by $7.1\%$ and operational cost by $3.0\%$ compared to the linear model, as it accurately captures GPU/TPU power dynamics and avoids over/underestimating CPN load. Meanwhile, communication constraints improve task success rate by $5.7\%$ and reduce average tardiness by $57.1\%$, as they prevent migration failures caused by bandwidth/latency bottlenecks. Moreover, the carbon-aware DRL reduces emissions by $28.3\%$ without significant economic loss, proving its ability to balance sustainability and cost. Furthermore, Two-stage synergy reduces operational cost by $8.4\%$ and RES curtailment rate by $55.6\%$, highlighting the value of real-time feedback to day-ahead planning. Therefore, we integrate the full framework of these modules to achieve optimal comprehensive performance.
V-G Analysis of Computational Burden
To verify the engineering feasibility of the TSCO framework, we quantify the computational time of each stage under a standard server environment (CPU: Intel Xeon 8375C 3.0GHz; GPU: NVIDIA A100 40GB; Memory: 256GB). The results are shown in Table IV.
TABLE IV: Computational Burden Data
Stage Metric Value (mins)
Day-Ahead SUC Average Convergence Iterations 27.8±2.3
Average Time per Iteration (min) 1.55±0.12
Average Total Convergence Time (min) 43.1±3.7
Master Problem Solving Time (min/iter) 0.32±0.05
Subproblems Solving Time (min/iter, parallel) 1.23±0.10
Real-Time Operation ED Average Solving Time (s/step) 1.2±0.1
DRL Inference Time (ms/task) 8.5±0.6
DRL Training Total Training Time (hours) 48.3±3.2
The day-ahead SUC stage has an average convergence time of $43.1$ minutes, which is far less than the $24$-hour planning cycle, leaving sufficient time for scenario updates and human review. The real-time ED solves in $1.2$ seconds per step, namely a $5$-minute interval. Meanwhile, DRL infers in $8.5$ milliseconds per task, meeting the requirements of real-time scheduling with second-level response. The offline DRL training takes $48.3$ hours, which is acceptable as it only needs to be performed once before deployment with incremental fine-tuning every $3$ months based on new data. These quantitative results confirm that the TSCO framework balances optimization accuracy and computational efficiency, making it suitable for practical engineering deployment.